### v2

@@ Abstract
Since optimization on Riemannian manifolds relies on the chosen metric, it is appealing to know that how the performance of a Riemannian optimization method varies with different metrics and how to exquisitely construct a metric such that a method can be accelerated. To this end, we propose a general framework for optimization problems on product manifolds endowed with a preconditioned metric, and we develop the Riemannian gradient descent and Riemannian conjugate gradient methods under this metric. Generally, the metric is constructed by an operator that aims to approximate the diagonal blocks of the Riemannian Hessian of the cost function. We propose three specific approaches to design the operator: exact block diagonal preconditioning, left and right preconditioning, and Gauss–Newton type preconditioning. In addition, we tailor new preconditioned metrics and adapt the proposed Riemannian methods to the canonical correlation analysis and the truncated singular value decomposition problems, and we adopt the Gauss–Newton type preconditioning to solve the tensor ring completion problem. Numerical results among these applications verify that a delicate metric does accelerate the Riemannian optimization methods.
@@ 1 Introduction
Recently, Riemannian optimization, designing algorithms based on the geometry of a Riemannian manifold $\mathcal{M}$, appears to be prosperous in many areas. Specifically, one can propose Riemannian optimization methods to solve problem Eq. 1, e.g., Riemannian gradient descent and Riemannian conjugate gradient methods. We refer to [2, 4] for a comprehensive overview.
@@
$\min\ f(\mathbf{x}):=-\mathbf{b}^{\mathsf{T}}\mathbf{x},\quad\mathrm{s.\,t.}\ \mathbf{x}\in\mathcal{M}_{\mathbf{B}}:=\{\mathbf{x}\in\mathbb{R}^{n}:\ \mathbf{x}^{\mathsf{T}}\mathbf{B}\mathbf{x}=1\},$
@@
$g_{\lambda,\mathbf{x}}:=\langle\xi,(\lambda\mathbf{I}_{n}+(1-\lambda)\mathbf{B})\eta\rangle\quad\text{for tangent vectors }\xi\text{ and }\eta,$
to the Riemannian gradient descent (RGD) method and the condition number of $\operatorname{Hess}_{g_{\lambda,\mathbf{x}}}\!f(\mathbf{x}^{*})$ in  Fig. 1, where $\lambda\in\mathbb{R}$ such that $\lambda\mathbf{I}_{n}+(1-\lambda)\mathbf{B}$ is positive definite. The left figure depicts the sequences generated by RGD under the Euclidean metric $g_{1,\mathbf{x}}(\xi,\eta)=\langle\xi,\eta\rangle$ and the scaled metric $g_{0,\mathbf{x}}(\xi,\eta)=\langle\xi,\mathbf{B}\eta\rangle$, and it shows that RGD under the metric $g_{0,\mathbf{x}}$ converges faster than the one under the Euclidean metric. Furthermore, the right figure confirms that the condition number varies with the metrics and $g_{0,\mathbf{x}}$ leads to the smallest condition number. The detailed computation can be found in Appendix A.
Figure 1: Left: sequences generated by the Riemannian gradient descent method under two metrics for $\mathbf{B}=\operatorname{diag}(2^{2},3^{2},1)$ and $\mathbf{b}=(1,1,1)$. Right: the condition number of $\operatorname{Hess}_{g_{\lambda,\mathbf{x}}}\!f(\mathbf{x}^{*})$ for $\lambda\in(-1/8,1]$. Blue marker: the Euclidean metric; green marker: the scaled metric.
@@
We propose three types of preconditioning approaches. The first is exact block diagonal preconditioning. Instead of approximating the full Riemannian Hessian, which can be computationally unfavorable in practice, we benefit from the block structure of $\operatorname{Hess}_{\mathrm{e}}\!f(x)$ and construct a new metric by exploiting the diagonal blocks; see an illustration in Fig. 2. Specifically, given $x=(x_{1},x_{2},\dots,x_{K})\in\mathcal{M}$, we can take advantage of the diagonal blocks $H_{kk}(x)$ of $\operatorname{Hess}_{\mathrm{e}}\!f(x)$ and construct a metric
@@
see Section 3.2. The Riemannian metric on $\mathcal{M}$ for both diagonal types is defined by the sum of the Riemannian metric on each component, i.e., $g_{x}(\xi,\eta)=\sum_{k=1}^{K}g_{x_{k}}^{k}(\xi_{k},\eta_{k})$. Thirdly, we propose Gauss–Newton type preconditioning technique for minimization of $\frac{1}{2}\|F(x)\|_{2}^{2}$ by constructing the metric
@@
By virtue of the new metric, we propose the Riemannian gradient descent and Riemannian conjugate gradient methods, and the condition number-related convergence results are presented. The preconditioned metric expands the scope of Riemannian preconditioning in [19] since it facilitates flexible choices of the operator ¯${\bar{\mathcal{H}}}(x)$. It is worth noting that exploiting more second-order information can improve the performance of Riemannian methods, but there is a trade-off between the increased cost brought by preconditioned metrics and the efficiency of preconditioned methods. The existing preconditioning methods that can be interpreted by the proposed framework are listed in Table 1.
Table 1: Existing and our works interpreted by preconditioned metrics. MC: matrix completion; TC: tensor completion; CP: canonical polyadic; TT: tensor train; TR: tensor ring.
@@ 2 Optimization on product manifolds
In this section, we provide basic tools in Riemannian geometry on product manifolds and develop the Riemannian gradient descent and Riemannian conjugate gradient methods for optimization on product manifolds; see [2, 4] for an overview. Metric-based and condition number-related convergence properties are presented.
@@ 2.1 Riemannian optimization on product manifolds
Assume that $\mathcal{M}$ is embedded in a Euclidean space $\mathcal{E}=\mathcal{E}_{1}\times\mathcal{E}_{2}\times\cdots\times\mathcal{E}_{K}$, which is called the ambient space. The tangent space of $\mathcal{M}$ at $x=(x_{1},x_{2},\dots,x_{K})$ is denoted by $\operatorname{T}_{x}\!\mathcal{M}=\operatorname{T}_{x_{1}}\!\mathcal{M}_{1}\times\operatorname{T}_{x_{2}}\!\mathcal{M}_{2}\times\cdots\times\operatorname{T}_{x_{K}}\!\mathcal{M}_{K}$, and a tangent vector is denoted by $\eta=(\eta_{1},\eta_{2},\dots,\eta_{K})$. Let each manifold $\mathcal{M}_{k}$ be endowed with a Riemannian metric $g^{k}$, and hence the Riemannian metric on the product manifold $\mathcal{M}$ can be defined by
@@ 3.1 Exact block diagonal preconditioning
Recall that a Riemannian metric on the product manifold $\mathcal{M}$ is defined by the sum of the metrics on each component, i.e., $g_{x}(\xi,\eta)=\sum_{k=1}^{K}g_{x_{k}}^{k}(\xi_{k},\eta_{k})$ for $\xi=(\xi_{1},\xi_{2},\dots,\xi_{K}),\eta=(\eta_{1},\eta_{2},\dots,\eta_{K})\in\operatorname{T}_{x}\!\mathcal{M}$. Consider the smooth extension $\bar{H}_{kk}(x):\operatorname{T}_{x_{k}}\!\mathcal{E}_{k}\to\operatorname{T}_{x_{k}}\!\mathcal{E}_{k}$ of the diagonal block $H_{kk}$ for $k=1,\dots,K$. If $\bar{H}_{11},\bar{H}_{22},\dots,\bar{H}_{KK}$ are positive definite on the ambient space $\mathcal{E}$, one can immediately adopt these blocks to construct the operator ${\bar{\mathcal{H}}}$ by
@@ 4.1 Left preconditioner
In our framework, it is equivalent that the operators in Eq. 5 are defined by ¯${\bar{\mathcal{H}}}_{1}(\mathbf{U},\mathbf{V})[\eta_{1}]=\Sigma_{xx}\eta_{1}$ and ¯${\bar{\mathcal{H}}}_{2}(\mathbf{U},\mathbf{V})[\eta_{2}]=\Sigma_{yy}\eta_{2}$, which have left preconditioning effect. The orthogonal projection with respect to $g$ of a vector ${\bar{\eta}}\in\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!\mathcal{E}\simeq\mathcal{E}$ onto $\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!\mathcal{M}$ is given by $\Pi_{g,(\mathbf{U},\mathbf{V})}({\bar{\eta}})=\left({\bar{\eta}}_{1}-\mathbf{U}\operatorname{sym}(\mathbf{U}^{\mathsf{T}}\Sigma_{xx}{\bar{\eta}}_{1}),{\bar{\eta}}_{2}-\mathbf{V}\operatorname{sym}(\mathbf{V}^{\mathsf{T}}\Sigma_{yy}{\bar{\eta}}_{2})\right)$, where $\mathcal{E}=\mathbb{R}^{d_{x}\times m}\times\mathbb{R}^{d_{y}\times m}$ is the ambient space of $\mathcal{M}$. Therefore, it follows from Eq. 4 that the Riemannian gradient reads
@@
A proof is given in [25, Theorem 5]. Nevertheless, we provide a slightly different proof for Proposition 4.1 in Appendix C for the sake of brevity, since the proofs of Propositions 4.4 and 5.1 stem from the details of the proof of Proposition 4.1. Note that Proposition 4.1 boils down to $\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))=(\sigma_{1}+\sigma_{2})/(\sigma_{1}-\sigma_{2})$ for $m=1$, which coincides with the result in [26, Lemma 4.1].
@@ 4.2 New left and right preconditioner
(10) $\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))=\frac{\max\{\underset{i,j\in[m],i\neq j}{\max}\frac{(\mu_{i}+\mu_{j})(\sigma_{i}+\sigma_{j})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta}},\frac{\mu_{1}(\sigma_{1}+\sigma_{m+1})}{\sqrt{\mu_{1}^{2}\sigma_{1}^{2}+\delta}}\}}{\min\{\underset{i,j\in[m],i\neq j}{\min}\frac{(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta}},\frac{\mu_{m}(\sigma_{m}-\sigma_{m+1})}{\sqrt{\mu_{m}^{2}\sigma_{m}^{2}+\delta}}\}}.$
@@
We aim to compute the Rayleigh quotient Eq. 4 under the proposed metric Eq. 7, and evaluate its upper and lower bounds. To this end, we firstly simplify the Riemannian Hessian $\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*})[\eta]$ as follows.
@@
$=\Pi_{\mathrm{new},(\mathbf{U}^{*},\mathbf{V}^{*})}($ $-\Sigma_{xx}^{-1}\Sigma_{xy}\eta_{2}\mathbf{N}(\mathbf{M}_{1,2}^{*})^{-1}+\eta_{1}\Sigma\mathbf{N}(\mathbf{M}_{1,2}^{*})^{-1}-\mathbf{U}^{*}\dot{\mathbf{S}}_{1}(\mathbf{M}_{1,2}^{*})^{-1},$
$-\Sigma_{yy}^{-1}\Sigma_{xy}^{\mathsf{T}}\eta_{1}\mathbf{N}(\mathbf{M}_{2,2}^{*})^{-1}+\eta_{2}\Sigma\mathbf{N}(\mathbf{M}_{2,2}^{*})^{-1}-\mathbf{V}^{*}\dot{\mathbf{S}}_{2}(\mathbf{M}_{2,2}^{*})^{-1}),$
where $\mathbf{M}_{1,2}^{*}=\mathbf{M}_{2,2}^{*}=(\Sigma^{2}\mathbf{N}^{2}+\delta\mathbf{I}_{m})^{1/2}$.
@@
Consequently, we can evaluate the upper and lower bounds of $q(\eta)$ in a similar fashion as Appendix C.
@@
Denote $\bar{v}_{ij}(\delta):=(\mu_{i}+\mu_{j})(\sigma_{i}+\sigma_{j})/(\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta})$ and $\underline{v}_{ij}(\delta):=(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j})/(\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta})$. It holds that
$\bar{v}_{ij}(0)>\bar{v}_{ik}(0)>\bar{v}_{jk}(0)\qquad\text{and}\qquad\underline{v}_{ij}(0)<\underline{v}_{ik}(0)<\underline{v}_{jk}(0)$
for all $1\leq i<j<k\leq m$ with $m\geq 3$.
@@
First, we prove that $\bar{v}_{ij}(0)>\bar{v}_{ik}(0)$. Since $\mu_{i}>\mu_{j}>\mu_{k}$ and $\sigma_{i}>\sigma_{j}>\sigma_{k}$, it holds that
@@
Therefore, $\bar{v}_{ij}(0)>\bar{v}_{ik}(0)$ holds and $\bar{v}_{ik}(0)>\bar{v}_{jk}(0)$ is also true. The result $\underline{v}_{ij}(0)<\underline{v}_{ik}(0)<\underline{v}_{jk}(0)$ is ready by using $\bar{v}_{ij}(0)+\underline{v}_{ij}(0)=2$, $\bar{v}_{ik}(0)+\underline{v}_{ik}(0)=2$ and $\bar{v}_{jk}(0)+\underline{v}_{jk}(0)=2$.
Then, we can prove that $\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))\leq\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ for $m\geq 2$. Note that $\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))=\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ for $m=1$ since the right preconditioners in Eq. 7 boil down to scalars that no longer have preconditioning effect.
@@
First, it follows from Lemma 4.6 and the continuity of $\bar{v}_{ij}$ and $\underline{v}_{ij}$ with respect to $\delta\in[0,\infty)$ that there exists a constant $\bar{\delta}>0$, such that
$\bar{v}_{ij}(\delta)>\bar{v}_{ik}(\delta)>\bar{v}_{jk}(\delta)\qquad\text{and}\qquad\underline{v}_{ij}(\delta)<\underline{v}_{ik}(\delta)<\underline{v}_{jk}(\delta)$
hold for all $1\leq i<j<k\leq m$ and $\delta\in(0,\bar{\delta})$. Therefore, we have $\bar{v}_{12}(\delta)=\max_{i,j\in[m],i\neq j}\bar{v}_{ij}(\delta)$ and $\underline{v}_{12}(\delta)=\min_{i,j\in[m],i\neq j}\underline{v}_{ij}(\delta)$. Subsequently, we observe that

### v3

@@ Abstract
Since optimization on Riemannian manifolds relies on the chosen metric, it is appealing to know that how the performance of a Riemannian optimization method varies with different metrics and how to exquisitely construct a metric such that a method can be accelerated. To this end, we propose a general framework for optimization problems on product manifolds endowed with a preconditioned metric, and we develop Riemannian methods under this metric. Generally, the metric is constructed by an operator that aims to approximate the diagonal blocks of the Riemannian Hessian of the cost function. We propose three specific approaches to design the operator: exact block diagonal preconditioning, left and right preconditioning, and Gauss–Newton type preconditioning. Specifically, we tailor new preconditioned metrics and adapt the proposed Riemannian methods to the canonical correlation analysis and the truncated singular value decomposition problems, which provably accelerate the Riemannian methods. Additionally, we adopt the Gauss–Newton type preconditioning to solve the tensor ring completion problem. Numerical results among these applications verify that a delicate metric does accelerate the Riemannian optimization methods.
@@ 1 Introduction
Riemannian optimization, designing algorithms based on the geometry of a Riemannian manifold $\mathcal{M}$, appears to be prosperous in many areas. One can propose Riemannian optimization methods to solve problem Eq. 1, e.g., Riemannian gradient descent and Riemannian conjugate gradient methods. We refer to [2, 4] for a comprehensive overview.
@@
$\min\ f(\mathbf{x}):=-\mathbf{b}^{\mathsf{T}}\mathbf{x},\quad\mathrm{s.\,t.}\ \mathbf{x}\in\mathcal{M}_{\mathbf{B}}:=\{\mathbf{x}\in\mathbb{R}^{n}:\ \mathbf{x}^{\mathsf{T}}\mathbf{B}\mathbf{x}=1\},=$
@@
$g_{\lambda,\mathbf{x}}(\xi,\eta):=\langle\xi,(\lambda\mathbf{I}_{n}+(1-\lambda)\mathbf{B})\eta\rangle\quad\text{for tangent vectors }\xi\text{ and }\eta,$
to the Riemannian gradient descent (RGD) method and the condition number of $\operatorname{Hess}_{g_{\lambda}}\!f(\mathbf{x}^{*})$ in  Fig. 1, where $\lambda\in\mathbb{R}$ such that $\lambda\mathbf{I}_{n}+(1-\lambda)\mathbf{B}$ is positive definite. The left figure depicts the sequences generated by RGD under the Euclidean metric $g_{1,\mathbf{x}}(\xi,\eta)=\langle\xi,\eta\rangle$ and the scaled metric $g_{0,\mathbf{x}}(\xi,\eta)=\langle\xi,\mathbf{B}\eta\rangle$, and it shows that RGD under the metric $g_{0}$ converges faster than the one under the Euclidean metric. Furthermore, the right figure confirms that the condition number varies with the metrics and $g_{0}$ leads to the smallest condition number. The detailed computation can be found in Appendix A.
Figure 1: Left: sequences generated by the Riemannian gradient descent method under two metrics for $\mathbf{B}=\operatorname{diag}(2^{2},3^{2},1)$ and $\mathbf{b}=(1,1,1)$. Right: the condition number of $\operatorname{Hess}_{g_{\lambda}}\!f(\mathbf{x}^{*})$ for $\lambda\in(-1/8,1]$. Blue marker: the Euclidean metric; green marker: the scaled metric.
@@
We propose three types of preconditioning approaches. The first type is exact block diagonal preconditioning. Instead of approximating the full Riemannian Hessian, which can be computationally unfavorable in practice, we benefit from the block structure of $\operatorname{Hess}_{\mathrm{e}}\!f(x)$ and construct a new metric by exploiting the diagonal blocks; see an illustration in Fig. 2. Specifically, given $x=(x_{1},x_{2},\dots,x_{K})\in\mathcal{M}$, we can take advantage of the diagonal blocks $H_{kk}(x)$ of $\operatorname{Hess}_{\mathrm{e}}\!f(x)$ and construct a metric
@@
see Section 3.2. The Riemannian metric on $\mathcal{M}$ for both diagonal types is defined by the sum of the Riemannian metric on each component, i.e.,
$g_{x}(\xi,\eta)=g_{x_{1}}^{1}(\xi_{1},\eta_{1})+g_{x_{2}}^{2}(\xi_{2},\eta_{2})+\cdots+g_{x_{K}}^{K}(\xi_{K},\eta_{K}).$
Thirdly, we propose Gauss–Newton type preconditioning technique for minimization of $\frac{1}{2}\|F(x)\|_{2}^{2}$ by constructing the metric
@@
By virtue of the new metric, we propose the Riemannian gradient descent and Riemannian conjugate gradient methods, and the condition number-related convergence results are developed. The preconditioned metric expands the scope of Riemannian preconditioning in [19] since it facilitates flexible choices of the operator ¯${\bar{\mathcal{H}}}(x)$. It is worth noting that exploiting more second-order information can improve the performance of Riemannian methods, but there is a trade-off between the increased cost brought by preconditioned metrics and the efficiency of preconditioned methods. The existing preconditioning methods that can be interpreted by the proposed framework are listed in Table 1.
Table 1: Existing and our works interpreted by preconditioned metrics. MC: matrix completion; TC: tensor completion; CP: canonical polyadic; TT: tensor train; TR: tensor ring; “*”: non-sigular matrices or tensors.
@@ 2 Optimization on product manifolds
In this section, we provide basic tools in Riemannian geometry on product manifolds and develop the Riemannian gradient descent and Riemannian conjugate gradient methods for optimization on product manifolds; see [2, 4] for an overview. Metric-based and condition number-related convergence properties are developed.
@@ 2.1 Riemannian optimization on product manifolds
Assume that $\mathcal{M}$ is embedded in a Euclidean space $\mathcal{E}=\mathcal{E}_{1}\times\mathcal{E}_{2}\times\cdots\times\mathcal{E}_{K}$, which is called the ambient space. The tangent space of $\mathcal{M}$ at $x=(x_{1},x_{2},\dots,x_{K})$ is denoted by $\operatorname{T}_{x}\!\mathcal{M}=\operatorname{T}_{x_{1}}\!\mathcal{M}_{1}\times\operatorname{T}_{x_{2}}\!\mathcal{M}_{2}\times\cdots\times\operatorname{T}_{x_{K}}\!\mathcal{M}_{K}$, and a tangent vector is denoted by $\eta=(\eta_{1},\eta_{2},\dots,\eta_{K})$. Let each manifold $\mathcal{M}_{k}$ be endowed with a Riemannian metric $g^{k}$. The Riemannian metric on the product manifold $\mathcal{M}$ can be defined by
@@ 3.1 Exact block diagonal preconditioning
Recall that a Riemannian metric on the product manifold $\mathcal{M}$ is defined by the sum of the metrics on each component, i.e., $g_{x}(\xi,\eta)=\sum_{k=1}^{K}g_{x_{k}}^{k}(\xi_{k},\eta_{k})$ for $\xi=(\xi_{1},\xi_{2},\dots,\xi_{K}),\eta=(\eta_{1},\eta_{2},\dots,\eta_{K})\in\operatorname{T}_{x}\!\mathcal{M}$. Consider the smooth extension $\bar{H}_{kk}(x):\operatorname{T}_{x_{k}}\!\mathcal{E}_{k}\to\operatorname{T}_{x_{k}}\!\mathcal{E}_{k}$ of the diagonal block $H_{kk}$ for $k=1,2,\dots,K$. If $\bar{H}_{11},\bar{H}_{22},\dots,\bar{H}_{KK}$ are positive definite on the ambient space $\mathcal{E}$, one can immediately adopt these blocks to construct the operator ${\bar{\mathcal{H}}}$ by
@@ 4.1 Left preconditioner
In our framework, it is equivalent that the operators in Eq. 5 are defined by ¯${\bar{\mathcal{H}}}_{1}(\mathbf{U},\mathbf{V})[\eta_{1}]=\Sigma_{xx}\eta_{1}$ and ¯${\bar{\mathcal{H}}}_{2}(\mathbf{U},\mathbf{V})[\eta_{2}]=\Sigma_{yy}\eta_{2}$, which have left preconditioning effect. The orthogonal projection with respect to $g$ of a vector ${\bar{\eta}}\in\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!\mathcal{E}\simeq\mathcal{E}$ onto $\operatorname{T}_{(\mathbf{U},\mathbf{V})}\!\mathcal{M}$ is given by $\Pi_{g,(\mathbf{U},\mathbf{V})}({\bar{\eta}})=\left({\bar{\eta}}_{1}-\mathbf{U}\operatorname{sym}(\mathbf{U}^{\mathsf{T}}\Sigma_{xx}{\bar{\eta}}_{1}),{\bar{\eta}}_{2}-\mathbf{V}\operatorname{sym}(\mathbf{V}^{\mathsf{T}}\Sigma_{yy}{\bar{\eta}}_{2})\right)$, where $\mathcal{E}=\mathbb{R}^{d_{x}\times m}\times\mathbb{R}^{d_{y}\times m}$ is the ambient space of $\mathcal{M}$. Therefore, it follows from Eq. 4 that the Riemannian gradient is
@@
A proof is given in [25, Theorem 5]. Nevertheless, we provide a slightly different proof for Proposition 4.1 in Appendix C to facilitate the proofs of Propositions 4.4 and 5.1. Specifically, the proof sketch of Propositions 4.1, 4.4, and 5.1 follows from the same procedure: 1) compute the Riemannian Hessian under the given metric $\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*})[\eta]$; 2) compute the Rayleigh quotient $q(\eta)$; 3) compute the maximum and minimum of Rayleigh quotient by taking the parametrization of tangent spaces into $q(\eta)$. Note that Proposition 4.1 boils down to $\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))=(\sigma_{1}+\sigma_{2})/(\sigma_{1}-\sigma_{2})$ for $m=1$, which coincides with the result in [26, Lemma 4.1].
@@ 4.2 New left and right preconditioner
(10) $\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))=\frac{\max\{\underset{i,j\in[m],i\neq j}{\max}\frac{(\mu_{i}+\mu_{j})(\sigma_{i}+\sigma_{j})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta}},\underset{i\in[m]}{\max}\frac{\mu_{i}(\sigma_{i}+\sigma_{m+1})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}}\}}{\min\{\underset{i,j\in[m],i\neq j}{\min}\frac{(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta}},\underset{i\in[m]}{\min}\frac{\mu_{i}(\sigma_{i}-\sigma_{m+1})}{\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}}\}}.$
@@
By following the same procedure as Proposition 4.1, we compute the Rayleigh quotient Eq. 4 under the proposed metric Eq. 7, and evaluate its upper and lower bounds. To this end, we firstly compute and simplify the Riemannian Hessian $\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*})[\eta]$ as follows.
@@
$=\Pi_{\mathrm{new},(\mathbf{U}^{*},\mathbf{V}^{*})}(-\Sigma_{xx}^{-1}\Sigma_{xy}\eta_{2}\mathbf{N}(\mathbf{M}_{1,2}^{*})^{-1}+\eta_{1}\Sigma\mathbf{N}(\mathbf{M}_{1,2}^{*})^{-1},$
$\qquad\qquad\qquad~{}~{}~{}-\!\Sigma_{yy}^{-1}\Sigma_{xy}^{\mathsf{T}}\eta_{1}\mathbf{N}(\mathbf{M}_{2,2}^{*})^{-1}+\eta_{2}\Sigma\mathbf{N}(\mathbf{M}_{2,2}^{*})^{-1}),$
where $\mathbf{M}_{1,2}^{*}=\mathbf{M}_{2,2}^{*}=(\Sigma^{2}\mathbf{N}^{2}+\delta\mathbf{I}_{m})^{1/2}$ are diagonal matrices, and we use the characterization of $(\operatorname{T}_{\mathbf{U}^{*}}\!\mathcal{M}_{1})^{\perp}$ and $(\operatorname{T}_{\mathbf{V}^{*}}\!\mathcal{M}_{2})^{\perp}$ in Eq. 1.
@@
We observe that only the denominator of $q(\eta)$ is different from the denominator $\langle\eta_{1},\Sigma_{xx}\eta_{1}\rangle+\langle\eta_{2},\Sigma_{yy}\eta_{2}\rangle$ in Eq. 2, and $\mathbf{M}_{1,2}^{*},\mathbf{M}_{2,2}^{*}$ are diagonal matrices. Consequently, we can evaluate the upper and lower bounds of $q(\eta)$ in a similar fashion as Appendix C.
@@
Denote 1) $\bar{v}_{ij}(\delta):=(\mu_{i}+\mu_{j})(\sigma_{i}+\sigma_{j})/(\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta})$ and $\bar{v}_{i,m+1}(\delta)=\mu_{i}(\sigma_{i}+\sigma_{m+1})/\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}$; 2) $\underline{v}_{ij}(\delta):=(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j})/(\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}+\sqrt{\mu_{j}^{2}\sigma_{j}^{2}+\delta})$ and $\underline{v}_{i,m+1}(\delta):=\mu_{i}(\sigma_{i}-\sigma_{m+1})/\sqrt{\mu_{i}^{2}\sigma_{i}^{2}+\delta}$ for $i,j\in[m]$. It holds that
$\bar{v}_{ij}(0)>\bar{v}_{ik}(0)\qquad\text{and}\qquad\underline{v}_{ij}(0)<\underline{v}_{ik}(0)$
for all $11\leq i<j<k\leq m+$ 1 with $m\geq 3$.
@@
We observe that $\bar{v}_{ij}(0)=(\mu_{i}+\mu_{j})(\sigma_{i}+\sigma_{j})/(\mu_{i}\sigma_{i}+\mu_{j}\sigma_{j})$ and $\underline{v}_{ij}(0)=(\mu_{i}-\mu_{j})(\sigma_{i}-\sigma_{j})/(\mu_{i}\sigma_{i}+\mu_{j}\sigma_{j})$ for $i\in[m]$ and $j\in[m+1]$ with $\mu_{m+1}=0$. First, we prove that $\bar{v}_{ij}(0)>\bar{v}_{ik}(0)$. Since $\mu_{i}>\mu_{j}>\mu_{k}$ and $\sigma_{i}>\sigma_{j}>\sigma_{k}$, it holds that
@@
Therefore, $\bar{v}_{ij}(0)>\bar{v}_{ik}(0)$ holds. The results $\underline{v}_{ij}(0)<\underline{v}_{ik}(0)$ is ready by using $\bar{v}_{ij}(0)+\underline{v}_{ij}(0)=2$ and $\bar{v}_{ik}(0)+\underline{v}_{ik}(0)=2$.
Then, it follows from Lemma 4.6 and the continuity of $\bar{v}_{ij}$ and $\underline{v}_{ij}$ with respect to $\delta\in[0,\infty)$ that there exists a constant $\bar{\delta}_{1}>0$, such that
$\bar{v}_{ij}(\delta)>\bar{v}_{ik}(\delta)\qquad\text{and}\qquad\underline{v}_{ij}(\delta)<\underline{v}_{ik}(\delta)$
hold for all $11\leq i<j<k\leq m+$ 1 and $\delta\in(0,\bar{\delta}_{1})$. Therefore, we can simplify the condition number
(11) $\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ $=\frac{\max_{i\in[m],j\in[m+1],i\neq j}\bar{v}_{ij}(\delta)}{\min_{i\in[m],j\in[m+1],i\neq j}\underline{v}_{ij}(\delta)}=\frac{\max_{i\in[m]}\bar{v}_{i,i+1}(\delta)}{\min_{i\in[m]}\underline{v}_{i,i+1}(\delta)}.$
We aim to prove that $\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))\leq\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ for $m\geq 2$. Note that $\kappa_{\mathrm{new}}(\operatorname{Hess}_{\mathrm{new}}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))=\kappa_{g}(\operatorname{Hess}_{g}\!f(\mathbf{U}^{*},\mathbf{V}^{*}))$ for $m=1$ since the right preconditioners in Eq. 7 boil down to scalars that no longer have preconditioning effect.
@@
If $\operatorname{arg\,max}_{i\in[m]}\bar{v}_{i,i+1}(0)=\{i^{*}\}$ for some $i^{*}\in[m]$, it follows from $\bar{v}_{i,i+1}(0)+\underline{v}_{i,i+1}(0)=2$ that $\operatorname{arg\,min}_{i\in[m]}\underline{v}_{i,i+1}(0)=\operatorname{arg\,max}_{i\in[m]}\bar{v}_{i,i+1}(0)=\{i^{*}\}$. Since $\bar{v}_{ij}(\delta)$ and $\underline{v}_{ij}(\delta)$ are continuous, there exists $\bar{\delta}>0$, such that $\{i^{*}\}=\operatorname{arg\,max}_{i\in[m]}\bar{v}_{i,i+1}(\delta)$ and $\{i^{*}\}=\operatorname{arg\,min}_{i\in[m]}\underline{v}_{i,i+1}(\delta)$ for all $\delta\in[0,\bar{\delta})$. Subsequently, we obtain that